OpenStudy (anonymous):

Find the length of a rectangle with area of 16 sq. cm. and width of 8 cm. 2 cm 8 cm 128 cm

OpenStudy (campbell_st):

the formula is A= lw so 16 = 8l or l = 16/2 l = 2 the length of the rectangle is 2 cm

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

Since area of rectangular is l*b so l*b=16 L=16/8 L=2